OpenStudy (eric_d):

Determine whether this equation defines y as a function of x Please explain @ganeshie8

OpenStudy (cggurumanjunath):

where is the function ?

OpenStudy (eric_d):

y^2=x-2 @ganeshie8

ganeshie8 (ganeshie8):

what exactly is needed for an equation to be qualified as function ?

ganeshie8 (ganeshie8):

heard of `vertical line test` before ?

ganeshie8 (ganeshie8):

For \(y\) to be a function of \(x\), every \(x\) value must point to exactly one \(y\) value : \(\large y^2 = x-2 \implies y = \pm \sqrt{x-2}\)

ganeshie8 (ganeshie8):

Can tell me the value of y, when x = 3 ?

OpenStudy (eric_d):

what exactly is needed for an equation to be qualified as function ? I'm not sure Vertical test line..! No +1, -1

OpenStudy (eric_d):

what exactly is needed for an equation to be qualified as function ? The value of y must be unique... but in terms of what?

ganeshie8 (ganeshie8):

That's a very good question ! Suppose you are `x`, do u have a girl friend ?

OpenStudy (eric_d):

yes

ganeshie8 (ganeshie8):

good, lets say your girl friend is `y`

OpenStudy (eric_d):

so, there must be one value only for y

ganeshie8 (ganeshie8):

here is the rule : if you(x) have `only one` girlfriend(y), then your girl friend(y) is qualified to be called as a function of x

ganeshie8 (ganeshie8):

if you have more than one girl friend, then none of your girl friends can be called a function ! Pardon for the bad example lol, but u get the idea i hope... a graph would clear things up I guess... let me show you the graph of given equation :)

OpenStudy (eric_d):

Np..!

OpenStudy (eric_d):

Understood

OpenStudy (eric_d):

Thanks

ganeshie8 (ganeshie8):

Look at the graph ^ for each x, there are two different possible y values - So y is not a function of x.

OpenStudy (eric_d):

Thanks for the example Another question: Sketch graph of function f(x)=x^2+4x+5 It can be expresses by completing the square =(x+2)^2+1 Thegraph cuts y-axis when x=0, y=5 Then..

OpenStudy (eric_d):

As x approach infinity, y approaches infinity Similarly when x approaches - infinity, y approaches infinity Why there's negative sign for ..

ganeshie8 (ganeshie8):

May be I'll answer the negative sign part after we sketch the curve : step 1 : plot the vertex

OpenStudy (eric_d):

 I know how to sketch.. I'm just understand the negative infinity part